OpenStudy (anonymous):

Write the following equation in standard form: 2 + 4(x - 3) + 2 = 4 + 5y. Show your work. Can someone explain how to solve this???

OpenStudy (anonymous):

okay :v, I'll try my best to answer and explain :3

OpenStudy (anonymous):

2 + 4(x - 3) + 2 = 4 + 5y : original 2 +2 +4 (x -3) =4 +5y : I just rewrite the whole thing :) looks easier 4 +4x -12 =4 +5y : distribute property 4 times x and 4 times -3 4x-8 =4+5y : subtraction +8 +8 : add 8 for both side 4x=12+5y : you can leave this is your answers or you can 4x-5y =12 :) you can't do anything more cuz we have 2 different values :) so this is your final answer :)

OpenStudy (anonymous):

Thank you so much!